On Q/{-1} define * by a*b=a+b+an for all a,b element of Q/{-1} show that (Q/{-1},*) is an abelian group
1
Expert's answer
2020-09-17T10:27:26-0400

"a\in Q/\{-1\}"

"b∈Q/\{−1\}"

"a*b=a+b+ab\in Q/\{-1\}"


"\therefore *, is"  a binary operation in"(Q/\{-1\}, *)"

Associativity: Suppose "a,b,c \in Q/\{-1\}"

"(a*b)*c=(a+b+ab)*c"

"=(a+b+ab)+c+(a+b+ab)c"

"=a+b+c+ab+bc+ca+abc"


So,

"a*(b*c)=a*(b+c+bc)"

"=a+(b+c+bc)+a(b+c+bc)"


"=a+(b+c+bc)+a(b+c+bc)"

"=a+b+c+ab+bc+ca+abc"

"(a*b)*c=a*(b*c)"

* is associative.


"Existence" "of" "Identity:"

"0\in Q/\{-1\}"

"is" the identity element.

Because "a\in Q/\{-1\}"

"0*a=0+a+0a=a"

"a*0=a+0+a0=a"

"Existence" of "Inverse" :

Every "a\in" "Q/\{-1\}"

For -a/a+1 "\in" Q/{-1}

"a\not=-1," which is inverse of a

Because

"\bigg(-a/(a+1) \bigg) *a" "="

"=\bigg(-a/(a+1)\bigg) +a+\bigg(-a/(a+1)\bigg) *a=0"


"a*\bigg(-a/(a+1)\bigg) ="

"=a+(-a)/(a+1)+a\bigg(-a/(a+1)\bigg) =0"


"Commutativity:-" "a, b \in Q/\{-1\}"

"a*b=a+b+ab"

"b*a=b+a+ba"


So,

"a*b=b*a"

Operation * is commutative.


The above discussion proves that the group is abelian.